One number is eight more than twice another. If their sum is decreased by nine, the result is fourteen. Find the numbers
Answer by nyc_function(2741)

n = 2m + 8
(n + m) - 9 = 14
=================

2m + 8 + m - 9 = 14

3m - 1 = 14

3m = 14 + 1

3m = 15

m = 15/3

m = 5

n = 2(5) + 8

n = 10 + 8

n = 18

The numbers are 5 and 18.